OpenStudy (anonymous):

Evaluate 4/7-7/15+1/3

OpenStudy (anonymous):

so what are you divideing here

OpenStudy (anonymous):

\[\frac{4}{7}-\frac{7}{15}+\frac{1}{3}\]?

OpenStudy (anonymous):

Ughhh I don't noe...I thought it was to add across

OpenStudy (anonymous):

Ya @satellite73

OpenStudy (anonymous):

I noe the answer but I don't get how to get that answer

OpenStudy (anonymous):

ohhh so you are doing fractions what confused me is the /  sorry my mouse went crazy lol

OpenStudy (anonymous):

Lol yeah its fractions

OpenStudy (anonymous):

common denominator is \(7\times 15=105\)

OpenStudy (anonymous):

\[\frac{4}{7}-\frac{7}{15}+\frac{1}{3}\] \[=\frac{4\times 15}{105}-\frac{7\times 7}{105}+\frac{1\times 35}{105}\]

OpenStudy (anonymous):

Wait thats confusing

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

No but answer is 2 over 3